Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono
ADS: E. Anaclerio

W unit

1) Final routing of WFS board cabling to accommodate the fabric cover has ben done.
2) The verification of the WFS board internal optical alignment activity has started today.
a) a measurement to quantify PSF and pupil displacement as a function of the rotation of the AGW frame. Data reduction is ongoing. No major problem has ben found in terms of displacements.
b) the geometry of the four pupil images has been measured and data reduction is ongoing.

AO control SW

1) Some work on AOS going on.
2) Installation tests on adsecdx. A few link problems have been investigated.
3) After the programmed poweroff the Clean Room switch lost configuration. Some time spent to diagnose. Jim Davis fixed this.
4) IDL license files set up to allow fast switch from the server based license to a standalone (Arcetri laptop file) license. A chapter about IDL added to documentation


LBT672a

1) Astatic lever has been connected to the unit.
2) Electronic cabinet has been placed in position.
3) Cooling lines have been connected to the electronic cabinet and tested.


Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback